# Exploring Variance Reduction in Importance Sampling for Efficient DNN Training

### v1

@@ 5.2 Application to automatic learning rate adjustment
6 Designing importance-score-estimation algorithm using variance estimators
@@ 6.1 Absolute metric for effectiveness of importance sampling
  • When $\mathcal{S}(W)$ is close to 00, the $W$ perform nearly as well as $W^{\ast}$.
@@
Therefore, if $10\leq\mathcal{S}(W)<$ 1, importance sampling effectively reduces the variance, with smaller $\mathcal{S}(W)$ indicating greater variance reduction. Conversely, if $\mathcal{S}(W)\geq 1$, importance sampling fails to reduce the variance compared with uniform sampling.
@@ 6.2 Gradient-norm estimation using moving statistics
Our Algorithm 1 estimates $W$ using moving statistics of $g_{i}^{t}$ during training. In Algorithm 1, the moving average of $g_{i}^{t}$ is maintained as the internal state $\hat{\mu}_{i}\ \left(i\in\llbracket M\rrbracket\right)$. At $t$, when $g_{i}^{t}$ is computed for $i$, the moving average is updated via UpdateStats. We assume that $g_{i}^{t}$ is computed only for data samples included in the minibatch at $t$, and UpdateStats is applied accordingly. Notably, when importance sampling is used, the intervals between iterations at which an $i$ appears in the minibatch become non-uniform. To address this, UpdateStats extends the exponential moving average (EMA) to account for non-uniform time intervals [6]. Furthermore, Algorithm 1 estimates the moving variance $\hat{\sigma}_{i}^{2}$ [7] alongside the moving average $\hat{\mu}_{i}$. The $W$ is then computed by summing these statistics in ComputeImportance. The inclusion of $\hat{\sigma}_{i}^{2}$ serves to prevent numerical instabilities, as discussed in [1, 11].
@@
[Hyper parameter] $\tau>0$
@@
8:     return $W=\left\{w_{i}\ \middle|\ w_{i}=\hat{\mu}_{i}+\sqrt{\hat{\sigma}_{i}^{2}},\quad\forall i\in\llbracket M\rrbracket\right\}$
@@
Hyperparameter
Hyperparameter $\tau$ in Algorithm 1 determines the extent to which past observations affect the moving statistics. For example, when $\tau$ is small, $\alpha$ computed in line 2 becomes small (closer to 00). Consequently, in line 4, the moving average $\hat{\mu}_{i}$ is updated to place greater emphasis on the current observation $g_{i}^{t}$ rather than past values. Conversely, when $\tau$ is large, $\alpha$ approaches $1$, causing the update to prioritize past values over the current observation. In the next section, we examine the effect of $\tau$ on $W$ estimation using $\mathcal{S}(W)$ and discuss strategies for determining an appropriate $\tau$.
@@ 6.3 Hyperparameter investigation using the proposed metric
A straightforward approach to determining $\tau$ is to vary it, conduct training, and selecting the $\tau$ that achieves the highest prediction accuracy on a validation dataset. However, metrics, such as prediction accuracy, can be affected by factors other than $\tau$, making it difficult to directly evaluate the quality of $\tau$ or the resulting $W$. To address this, we investigated the impact of $\tau$ on the effectiveness of importance sampling using $\mathcal{S}(W)$.
@@ 6.3.2 Dynamic $\tau$
On the basis of the above results, we propose varying $\tau$ during training instead of using a fixed value. Specifically, we set $\tau$ to the current $t$ by inserting $\tau\leftarrow t$ just before line 2 in Algorithm 1. Hereafter, we refer to this method as Linear-$\tau$, which notably eliminates the need for hyperparameter tuning in Algorithm 1.
The training results for CINIC-10 using Linear-$\tau$ are shown in Figure 1(a) (labeled as ”Linear” in the legend). Linear-$\tau$ consistently achieved small $\mathcal{S}(W)$ throughout the training process. These results suggest that in the early stages of training, importance sampling is more efficient when $w_{i}$ adapts to the most recent $g_{i}^{t}$. However, as training progresses, estimating $w_{i}$ on the basis of a longer-term history of $g_{i}^{t}$ leads to more stable importance sampling. The training results for FMNIST using Linear-$\tau$ are shown in Figure 1(b). While $\mathcal{S}(W)$ was slightly larger compared with when $\tau$ was fixed at 1000, it still maintained relatively small $\mathcal{S}(W)$ throughout the training process. Due to $\mathcal{S}(W)$, which efficiently assesses the effectiveness of importance sampling across the training process, a more refined and effective design of $\tau$ was achieved.
@@ 7.1 Datasets
The following three datasets were used in the evaluation: FMNIST [31] consists of $28\times 28$ grayscale images of clothing items labeled into ten classes. It contains 60000 images for training and 10000 images for testing. CINIC-10 [5] is a dataset of $32\times 32$ RGB images featuring objects such as animals and vehicles, labeled into ten classes. It contains 90000 images for training and 90000 images for testing. The ChestX-ray14 dataset [30] contains anonymized $1024\times 1024$ grayscale chest X-ray images. Labels for 14 types of diseases (e.g., pneumonia) are assigned to each image by processing radiologist reports using natural language processing methods. Since a single X-ray image may be associated with multiple diseases described in the corresponding report, the dataset uses multi-label annotations. Specifically, each image is assigned 14 binary labels, where a value of 1 indicates the presence of a particular disease and 0 indicates its absence. Images with all 14 labels set to 0 correspond to ”No Finding” (absence of any conditions). It contains 86524 images for training and 25596 images for testing.
@@ 7.2 Compared methods
  • RAIS [11]: In RAIS, $W$ is estimated using a robust regression model based on the per-sample gradient norm, and training is conducted using importance sampling with this $W$. The learning rate is also adjusted in accordance with the expected squared norm of the gradient. RAIS also introduces the ”effective iteration number,” a virtual iteration count used to adjust the learning rate scheduling accordingly.
@@
  • Self-paced learning [16]: Self-paced learning, a variant of curriculum learning [3], automatically estimates the difficulty of each data point and selects training data accordingly. This method introduces a hyperparameter $K$ such that, at each training iteration, only data points with loss values below $1/K$ are used for training. As training progresses and more data points achieve smaller loss values, the proportion of data used for training increases. We use the method proposed by [16] to adjust $K$. It is important to note that self-paced learning was originally designed for batch learning and cannot be directly applied to minibatch-based DNN training. To address this, we use a straightforward approach: within each minibatch, only data points with loss values below $1/K$ are used for training.
@@
For RAIS, the robust regression model was reused from the authors’ implementation [10]. All other methods were implemented from scratch. The hyperparameters for RAIS, such as the variables used in regression, were directly adopted from the authors’ implementation. For Presampling-IS, the large initial minibatch size was set to $N_{\mathrm{large}}=1024$. With Self-paced learning, the initial value of $K$ was set such that 60% of the training data were used at the beginning, and $\mu=1.2$ was used to increase $K$ during training. Note that the application of the Confidence and ConfVar methods to ChestX-ray14 is not straightforward, as it contains multi-labeled data. Therefore, these methods were excluded from evaluation on ChestX-ray14 in the experiments.

### v2

@@ 5.2 Application to automatic learning rate adjustment
It should be noted that the optimal learning rate scaling rule when changing the minibatch size depends on the choice of optimizer. The above definition of $\epsilon_{\mathrm{ems}}$ assumes the use of SGD as the optimizer. For the Adam optimizer, for example, it has been both empirically [26] and theoretically [22] suggested that, especially when the minibatch size is relatively small, the learning rate should be scaled by the square root of the minibatch size ratio. Therefore, when using the Adam optimizer, the learning rate can be adjusted based on $N_{\mathrm{ems}}$ by setting $\epsilon_{\mathrm{ems}}^{\mathrm{adam}}:=\sqrt{N_{\mathrm{ems}}/N}~\epsilon$ instead of using Eq. 8.
6 Designing importance-weight-estimation algorithm using variance estimators
@@ 6.1 Absolute metric for effectiveness of importance sampling
  • When $\mathcal{S}(W)$ is close to 0, the $W$ perform nearly as well as $W^{\ast}$.
@@
Therefore, if $0\leq\mathcal{S}(W)<1$, importance sampling effectively reduces the variance, with smaller $\mathcal{S}(W)$ indicating greater variance reduction. Conversely, if $\mathcal{S}(W)\geq 1$, importance sampling fails to reduce the variance compared with uniform sampling.
@@ 6.2 Gradient-norm estimation using moving statistics
Our Algorithm 1 estimates $W$ using moving statistics of $g_{i}^{t}$ during training. In Algorithm 1, the moving average of $g_{i}^{t}$ is maintained as the internal state $\hat{\mu}_{i}\ \left(i\in\llbracket M\rrbracket\right)$. At $t$, when $g_{i}^{t}$ is computed for $i$, the moving average is updated via UpdateStats. We assume that $g_{i}^{t}$ is computed only for data samples included in the minibatch at $t$, and UpdateStats is applied accordingly. Notably, when importance sampling is used, the intervals between iterations at which an $i$ appears in the minibatch become non-uniform. To address this, UpdateStats extends the exponential moving average (EMA) to account for non-uniform time intervals [7]. Furthermore, Algorithm 1 estimates the moving variance $\hat{\sigma}_{i}^{2}$ [8] alongside the moving average $\hat{\mu}_{i}$. The $W$ is then computed by summing these statistics in ComputeImportance. The inclusion of $\hat{\sigma}_{i}^{2}$ serves to prevent numerical instabilities, as discussed in [1, 13]. To prevent the sampling probabilities derived from $W$ from becoming excessively extreme, we do not directly use those computed by Eq. 1. Instead, we employ the ComputeAdjustedProbabilities procedure in Algorithm 1 to obtain a flattened distribution. This adjustment ensures that the expected maximum number of duplicates per minibatch does not exceed a specified hyperparameter $\kappa$, which governs the degree of flattening. In our experiments, we set $\kappa=1$.
@@
[Hyper parameter] $\tau(>0)$: time constant for exponential decay in the moving statistics, $N$: minibatch size, $\kappa$: expected maximum duplicates per minibatch
@@
8:  $w_{i}\leftarrow\hat{\mu}_{i}+\sqrt{\hat{\sigma}_{i}^{2}},\quad\forall i\in\llbracket M\rrbracket$
9:  return $W=\left(w_{1},\ldots,w_{M}\right)^{\top}$
10:function ComputeAdjustedProbabilities($W$)
11:  $p_{i}\leftarrow w_{i}/\sum_{i^{\prime}\in\llbracket M\rrbracket}w_{i^{\prime}},\quad\forall i\in\llbracket M\rrbracket$
12:  while $\max_{i}(p_{i})\times N>\kappa$ do
13:   $p_{i}\leftarrow\sqrt{p_{i}}/\sum_{i^{\prime}\in\llbracket M\rrbracket}\sqrt{p_{i^{\prime}}},\quad\forall i\in\llbracket M\rrbracket$
14:  return $\left(p_{1},\ldots,p_{M}\right)^{\top}$
@@
Hyperparameter for moving statistics
Hyperparameter $\tau$ in Algorithm 1 determines the extent to which past observations affect the moving statistics. For example, when $\tau$ is small, $\alpha$ computed in line 2 becomes small (closer to 0). Consequently, in line 4, the moving average $\hat{\mu}_{i}$ is updated to place greater emphasis on the current observation $g_{i}^{t}$ rather than past values. Conversely, when $\tau$ is large, $\alpha$ approaches $1$, causing the update to prioritize past values over the current observation. In the next section, we examine the effect of $\tau$ on $W$ estimation using $\mathcal{S}(W)$ and discuss strategies for determining an appropriate $\tau$.
@@ 6.3 Hyperparameter investigation using the proposed metric
A straightforward approach to determining $\tau$ is to vary it, conduct training, and selecting the $\tau$ that achieves the highest prediction accuracy on a validation dataset. However, metrics, such as prediction accuracy, can be affected by factors other than $\tau$, making it difficult to directly evaluate the quality of $\tau$ or the resulting $W$. To address this, we investigated the impact of $\tau$ on the effectiveness of importance sampling using $\mathcal{S}(W)$, which is estimated during training based on per-sample gradients.
@@ 6.3.2 Dynamic $\tau$
On the basis of the above results, we propose varying $\tau$ during training instead of using a fixed value. Specifically, we set $\tau$ to the current $t$ by inserting $\tau\leftarrow t$ just before line 2 in Algorithm 1. Hereafter, we refer to this method as Linear-$\tau$, which notably eliminates the need to tune the hyperparameter $\tau$ in Algorithm 1.
The training results for CINIC-10 using Linear-$\tau$ are shown in Figure 1(a) (labeled as “Linear” in the legend). Linear-$\tau$ consistently achieved small $\mathcal{S}(W)$ throughout the training process. These results suggest that in the early stages of training, importance sampling is more efficient when $w_{i}$ adapts to the most recent $g_{i}^{t}$. However, as training progresses, estimating $w_{i}$ on the basis of a longer-term history of $g_{i}^{t}$ leads to more stable importance sampling. The training results for FMNIST using Linear-$\tau$ are shown in Figure 1(b). While $\mathcal{S}(W)$ was slightly larger compared with when $\tau$ was fixed at 1000, it still maintained relatively small $\mathcal{S}(W)$ throughout the training process. Due to $\mathcal{S}(W)$, which efficiently assesses the effectiveness of importance sampling across the training process, a more refined and effective design of $\tau$ was achieved. As a reference, the efficiency score $\mathcal{S}(W)$ obtained using Linear-$\tau$, along with the corresponding estimated trace of gradient variances used in its computation, is provided in subsection C.1.
@@
7 Efficient per-sample gradient approximation via logit gradients
The formula in Proposition 4.1, along with the derived quantities $N_{\mathrm{ems}}$ and $\mathcal{S}(W)$, requires the computation of the loss gradient $\nabla_{\theta}\mathcal{L}(\theta;i)$ for each data point. However, calculating the loss gradient with respect to $\theta$ introduces significant computational overhead. To address this issue, methods have been proposed to use the gradient with respect to the output of the final layer of the DNN (referred to as the logits) instead of the gradient with respect to $\theta$ [13, 16]. Compared with computing the gradient for $\theta$, the gradient with respect to the logits can be obtained at a significantly lower computational cost.
By defining the logits as $z=f_{\theta}(x)$, the training loss function can be written as $\ell(z,y)$. We then introduce the auxiliary function $\tilde{\mathcal{L}}(z;i):=\ell(z,y^{(i)})$. Instead of computing the parameter gradient $\nabla_{\theta}\mathcal{L}(\theta;i)$, we first evaluate the logit gradient $\nabla_{z}\tilde{\mathcal{L}}(z;i):=\left.\nabla_{z}\ell(z,y^{(i)})\right|_{z=f_{\theta}(x^{(i)})}$, and substitute $\nabla_{z}\tilde{\mathcal{L}}(z;i)$ for $\nabla_{\theta}\mathcal{L}(\theta;i)$ in the formula in Proposition 4.1, resulting in the following expressions for the trace of the variance-covariance matrices of the gradient with respect to the logits:
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{z}\tilde{\mathcal{L}}(z)\right]\right)$ $=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[\left\|r(i;W)\nabla_{z}\tilde{\mathcal{L}}(z;i)\right\|^{2}\right]-\left\|\mu_{z}\right\|^{2},$ (10)
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{unif}}\left[\nabla_{z}\tilde{\mathcal{L}}(z)\right]\right)$ $=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\left\|\nabla_{z}\tilde{\mathcal{L}}(z;i)\right\|^{2}\right]-\left\|\mu_{z}\right\|^{2},$ (11)
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W^{\ast})}\left[\nabla_{z}\tilde{\mathcal{L}}(z)\right]\right)$ $=\left(\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\left\|\nabla_{z}\tilde{\mathcal{L}}(z;i)\right\|\right]\right)^{2}-\left\|\mu_{z}\right\|^{2},$ (12)
where $\mu_{z}$ is given by
$\mu_{z}$ $=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\nabla_{z}\tilde{\mathcal{L}}(z;i)\right].$
Note that, for example, in an $N_{\mathrm{c}}$-class classification problem, the dimension of $z$ is typically $N_{\mathrm{c}}$, and thus the dimension of $\nabla_{z}\tilde{\mathcal{L}}(z;i)$ is also $N_{\mathrm{c}}$. In contrast, the dimension of $\nabla_{\theta}\mathcal{L}(\theta;i)$ is $K$, corresponding to the number of trainable parameters. Since $N_{\mathrm{c}}\ll K$ in most cases, the two gradients differ in size, and their norms can vary substantially. However, as demonstrated by previous studies [13] and by our experiments in subsection 9.6, the norms of $\nabla_{z}\tilde{\mathcal{L}}(z;i)$ and $\nabla_{\theta}\mathcal{L}(\theta;i)$ exhibit a strong proportional relationship across different values of $i$. A similar trend is observed, for example, between $\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ and $\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{z}\tilde{\mathcal{L}}(z)\right]\right)$. Although the magnitudes of these quantities may differ significantly, our experiments (see subsection 9.6) empirically confirm a strong linear relationship between them.
While the estimated trace of the gradient variance based on the logits exhibits only a linear relationship with that computed from the parameter gradient, our experiments confirmed that the derived quantities $N_{\mathrm{ems}}$ and $S(W)$ can be estimated with a reasonable degree of absolute accuracy. This is attributed to the fact that $N_{\mathrm{ems}}$ and $S(W)$ are defined based on the ratio of variance traces and the ratio of differences between variance traces, respectively. The estimation errors of $N_{\mathrm{ems}}$ and $S(W)$ were also empirically evaluated in subsection 9.6.
Throughout the experiments in section 9, as well as the preliminary experiments in subsection 6.3.1, variance and score estimations were performed using gradients with respect to the logits as an approximation, except in subsection 9.6, which discusses the approximation accuracy of estimations based on the logit gradient. Note that $g_{i}^{t}$ in Algorithm 1 was also approximated using the logit gradients. Since the relationship $p_{\mathrm{is}}(i;W)=p_{\mathrm{is}}(i;\gamma W)$ holds for any $\gamma>0$, using logit-gradient norms in Algorithm 1 provides a reasonable basis for estimating $W$, assuming that the norm of the logit gradient is proportional to that of the parameter gradient.
8 Overall training procedure of EMAIS
The complete procedure for EMAIS training, incorporating the logit-based gradient approximation, is presented in Algorithm 2. It is important to note that, in Algorithm 2, both $\nabla_{z}\tilde{\mathcal{L}}(z;i_{k})$ and $\mathcal{L}(\theta;i_{k})$ are required for each data index $i_{k}$. However, in practice, these quantities can be computed simultaneously in a single forward pass rather than evaluated separately: we first compute $z^{(i_{k})}=f_{\theta}(x^{(i_{k})})$, then obtain $\nabla_{z}\tilde{\mathcal{L}}(z;i_{k})$ as $\left.\nabla_{z}\ell(z,y)\right|_{z=z^{(i_{k})}}$ and $\mathcal{L}(\theta;i_{k})$ as $\ell(z^{(i_{k})},y^{(i_{k})})$.
Algorithm 2 EMAIS training with logit-based gradient approximation
[…]
2:for $t=1$ to $t^{\prime}$ do $\triangleright$ Initialize EMAIS internal statistics
3:  Sample indices $\{i_{k}\}_{k=1}^{N}\sim p_{\mathrm{unif}}(i)$ without replacement
4:  for each $i_{k}$ do
5:   $\textsc{UpdateStats}\left(t,\left\|\nabla_{z}\tilde{\mathcal{L}}(z;i_{k})\right\|\right)$ $\triangleright$ See Algorithm 1
6:  $\hat{g}\leftarrow\nabla_{\theta}\left(\frac{1}{N}\sum_{k=1}^{N}\mathcal{L}(\theta;i_{k})\right)$
7:  $\theta\leftarrow\theta-\epsilon_{t}\hat{g}$
8:for $t=t^{\prime}+1$ to $T$ do
9:  $W\leftarrow\textsc{ComputeImportance}()$ $\triangleright$ See Algorithm 1
10:  $\left\{\tilde{p}_{i}\right\}_{i=1}^{M}\leftarrow\textsc{ComputeAdjustedProbabilities}(W)$ $\triangleright$ See Algorithm 1
11:  Sample indices $\{i_{k}\}_{k=1}^{N}$ according to $\left\{\tilde{p}_{i}\right\}_{i=1}^{M}$ with replacement
12:  for each $i_{k}$ do
13:   $g^{t}_{i_{k}}\leftarrow\nabla_{z}\tilde{\mathcal{L}}(z;\,i_{k})$
14:   $\textsc{UpdateStats}(t,\left\|g^{t}_{i_{k}}\right\|)$ $\triangleright$ See Algorithm 1
15:  Estimate $\phi_{\mathrm{is}}\approx\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{z}\tilde{\mathcal{L}}(z)\right]\right)$, $\phi_{\mathrm{unif}}\approx\operatorname{tr}\left(\mathbb{V}_{\mathrm{unif}}\left[\nabla_{z}\tilde{\mathcal{L}}(z)\right]\right)$, and $\phi_{\mathrm{ideal}}\approx\operatorname{tr}\left(\mathbb{V}_{\mathrm{is}(W^{\ast})}\left[\nabla_{z}\tilde{\mathcal{L}}(z)\right]\right)$ using Eqs. 10, 11 and 12, with $\left\{g^{t}_{i_{k}}\right\}_{k=1}^{N}$ and $r(i_{k},W)=(1/M)/\tilde{p}_{i_{k}}$. $\triangleright$ See also Proposition 4.1 and Figure 8
16:  $N_{\mathrm{ems}}\leftarrow\frac{\phi_{\mathrm{unif}}}{\phi_{\mathrm{is}}}N$
17:  $\epsilon_{\mathrm{ems}}\leftarrow\frac{N_{\mathrm{ems}}}{N}\epsilon_{t}$
18:  (Optional) $\mathcal{S}(W)\leftarrow\frac{\phi_{\mathrm{is}}-\phi_{\mathrm{ideal}}}{\phi_{\mathrm{unif}}-\phi_{\mathrm{ideal}}}$
19:  $\hat{g}\leftarrow\nabla_{\theta}\left(\frac{1}{N}\sum_{k=1}^{N}r(i_{k};W)\mathcal{L}(\theta;i_{k})\right)$
20:  $\theta\leftarrow\theta-\epsilon_{\mathrm{ems}}\hat{g}$
@@ 9.1 Datasets
The following three datasets were used in the evaluation: FMNIST [36] consists of $28\times 28$ grayscale images of clothing items labeled into ten classes. It contains 60000 images for training and 10000 images for testing. CINIC-10 [5] is a dataset of $32\times 32$ RGB images featuring objects such as animals and vehicles, labeled into ten classes. It contains 90000 images for training and 90000 images for testing. The ChestX-ray14 dataset [35] contains anonymized $1024\times 1024$ grayscale chest X-ray images. Labels for 14 types of diseases (e.g., pneumonia) are assigned to each image by processing radiologist reports using natural language processing methods. Since a single X-ray image may be associated with multiple diseases described in the corresponding report, the dataset uses multi-label annotations. Specifically, each image is assigned 14 binary labels, where a value of 1 indicates the presence of a particular disease and 0 indicates its absence. Images with all 14 labels set to 0 correspond to “No Finding” (absence of any conditions). It contains 86524 images for training and 25596 images for testing.
@@ 9.2 Compared methods
  • RAIS [13]: In RAIS, $W$ is estimated using a robust regression model based on the per-sample gradient norm, and training is conducted using importance sampling with this $W$. The learning rate is also adjusted in accordance with the expected squared norm of the gradient. RAIS also introduces the “effective iteration number,” a virtual iteration count used to adjust the learning rate scheduling accordingly.
@@
  • Self-paced learning [19]: Self-paced learning, a variant of curriculum learning [3], automatically estimates the difficulty of each data point and selects training data accordingly. This method introduces a hyperparameter $K_{\mathrm{sp}}$ such that, at each training iteration, only data points with loss values below $1/K_{\mathrm{sp}}$ are used for training. As training progresses and more data points achieve smaller loss values, the proportion of data used for training increases. We use the method proposed by [19] to adjust $K_{\mathrm{sp}}$. It is important to note that self-paced learning was originally designed for batch learning and cannot be directly applied to minibatch-based DNN training. To address this, we use a straightforward approach: within each minibatch, only data points with loss values below $1/K_{\mathrm{sp}}$ are used for training.
@@
For RAIS, the robust regression model was reused from the authors’ implementation [12]. All other methods were implemented from scratch. The hyperparameters for RAIS, such as the variables used in regression, were directly adopted from the authors’ implementation. For Presampling-IS, the large initial minibatch size was set to $N_{\mathrm{large}}=1024$. With Self-paced learning, the initial value of $K_{\mathrm{sp}}$ was set such that 60% of the training data were used at the beginning, and $\mu=1.2$ was used to increase $K_{\mathrm{sp}}$ during training. Note that the application of the Confidence and ConfVar methods to ChestX-ray14 is not straightforward, as it contains multi-labeled data. Therefore, these methods were excluded from evaluation on ChestX-ray14 in the experiments.